Bank X has Sh. 100M of assets and Sh. 10M of Equity bank. Bank Y has Sh. 100M in assets and Sh. 4Min equity. Calculate the Equity Multiplier of each bank. If the two banks have 1%, return on assets; calculate the return on equity of bank X and Y and comment. (10 Marks)
1
Expert's answer
2019-02-21T10:49:08-0500

Bank X has Sh. 100M of assets and Sh. 10M of Equity bank. Bank Y has Sh. 100M in assets and Sh. 4Min equity. The Equity Multipliers of each bank are:

EMx = Assets/Equity = 100M/10M = 10,

EMy = 100M/4M = 25.

If the two banks have 1%, return on assets, then the return on equity of bank X and Y are:

ROEx = ROA*EM = 0.01*10 = 0.1,

ROEy = 0.01*25 = 0.25.